Decision on transfer pricing

Materials used by the Instrument Division of XPort Industries are currently purchased from outside suppliers at a cost of $400 per unit. However, the same materials are available from the Components Division. The Components Division has unused capacity and can produce the materials needed by the Instrument Division at a variable cost of $332 per unit.

Assume that a transfer price of $380 has been established and that 22,600 units of materials are transferred, with no reduction in the Components Division’s current sales.

a. How much would XPort Industries’ total income from operations increase?
$

b. How much would the Instrument Division’s income from operations increase?
$

c. How much would the Components Division’s income from operations increase?
$
Answer #1

Answer a

XPort Industries total operating income increased by $ 1,536,800

Explanation:

Net benefit = (400-332)*22,600

= 1,536,800

Answer b

Instrument Division's operating income increased by $ 452,000.

Explanation:

Net benefit = (400-380)*22,600

= 452,000

Answer c

Components Division's operating income increased by $ 1,084,800.

Explanation:

Net benefit = (380-332)*22,600

= 1,084,800

In case of any doubt, please comment.